OpenStudy (anonymous):

Solve for W. -5(w+3)=-3w-9

OpenStudy (anonymous):

multiply the brackets out

OpenStudy (anonymous):

-5w-15=-3w-9 -15+9=-3w+5w -6=2w w=-3